The energy flux of sunlight reaching the surface of the earth is 1.388 × 103 W/m2. How many photons (nearly) per square metre are incident on the Earth per second? Assume that the photons in the sunlight have an average wavelength of 550 nm.

Grade:11

1 Answers

Vikas TU
3 years ago
Energy flux of sunlight reaching the surface of earth, Φ = 1.388 × 10^3 W/m^2
Hence, power of sunlight per square metre, P = 1.388 × 10^3 W
Speed of light, c = 3 × 10^8 m/s
Planck’s constant, h = 6.626 × 10^−34 Js
Average wavelength of photons present in sunlight, λ = 550 nm = 550 x 10^-9 m
Number of photons per square metre incident on earth per second = n Hence, the equation for power can be written as: P= nE
n = P/E = Pλ / hc = 1.388 x 10^3 x 550 x 10^-9 / 6.626 x 10^-34 x 3 x 10^8 = 3.84 x 10^21 photons/m2/s
Therefore, every second, 3.84 x 10^21 photons are incident per square metre on earth.